# When is Multicalibration Post-Processing Necessary?

### v1

@@ Abstract
Calibration is a well-studied property of predictors which guarantees meaningful uncertainty estimates. Multicalibration is a related notion — originating in algorithmic fairness — which requires predictors to be simultaneously calibrated over a potentially complex and overlapping collection of protected subpopulations (such as groups defined by ethnicity, race, or income). We conduct the first comprehensive study evaluating the usefulness of multicalibration post-processing across a broad set of tabular, image, and language datasets for models spanning from simple decision trees to 90 million parameter fine-tuned LLMs. Our findings can be summarized as follows: (1) models which are calibrated out of the box tend to be relatively multicalibrated without any additional post-processing; (2) multicalibration post-processing can help inherently uncalibrated models; and (3) traditional calibration measures may sometimes provide multicalibration implicitly. More generally, we also distill many independent observations which may be useful for practical and effective applications of multicalibration post-processing in real-world contexts.
@@ 1 Introduction
Although there is a host of theoretical results surrounding multicalibration and related notions, there is little systematic empirical study of the latent multicalibration error of popular machine learning models, the effectiveness of multicalibration post-processing algorithms, or even best practices for practitioners who wish to apply ideas and algorithms from the multicalibration literature. In particular, theoretical results are often concerned with multicalibration towards subgroups defined by potentially infinite hypothesis classes (Hébert-Johnson et al., , 2018; Haghtalab et al., , 2023), whereas fairness practitioners may only consider performance over a finite number of protected subgroups of interest, typically defined by attributes and meta-data normatively deemed as important (race, sex, ethnicity, etc. Chen et al., (2023)). Furthermore, most existing works applying multicalibration in practical settings only focus on one-off datasets or examples, and do not validate the algorithm across a variety of datasets and models or with realistic finite sample restrictions (Liu et al., , 2019; Barda et al., , 2020; La Cava et al., , 2023).
To address these, we consider a “realistic” setup where a practitioner only has a finite amount of data, and must choose how to partition this data between learning and post-processing in order to achieve a suitable accuracy and multicalibration error rate over a finite set of subgroups. This allows us to investigate many important questions pertaining to the practical usage of multicalibration concepts and algorithms, which, to the best of our knowledge, have not been systematically previously considered by the theoretical or practical communities. For example, we can investigate the effectiveness of multicalibration post-processing algorithms and hyperparameter choices, as well as the latent multicalibration properties of popular machine learning models at a large scale.
More broadly, we initiate a systematic empirical study of multicalibration with the goal of answering the following two salient questions:
@@
The conventional wisdom is that multicalibration is something that is not naturally achieved by ML algorithms—this is precisely why many in the community have focused on creating post-processing algorithms which do achieve it (see e.g. Hébert-Johnson et al., (2018); Gopalan et al., 2022b , and Section 1.2). However, recent theoretical results suggest that multicalibration may in fact be an inevitable consequence of certain empirical risk minimization (ERM) methods with proper losses (Błasiok et al., , 2023; Liu et al., , 2019). This apparent conflict between conventional wisdom and recent results has not been tested in practice. We propose studying Question 1 since we believe that the current state of multicalibration in ML models should be systematically studied to better understand the implications for modern learning setups involving large models and fine-tuning. Question 2 is complementary and focused on investigating the effectiveness of current multicalibration algorithms on real datasets and illuminating challenges which can guide the development of future algorithms. A partial answer to one or both of these questions could help practitioners concerned about fairness understand when they should or should not expect multicalibration algorithms to improve fairness.
Refer to captionRefer to caption
Figure 1: Test accuracy vs. maximum group-wise calibration error (smECE) averaged over five train/validation splits for MLP trained on CreditDefault and MEPS. Each point corresponds to the performance of the multicalibration post-processing algorithm HKRR (Hébert-Johnson et al., , 2018) or HJZ (Haghtalab et al., , 2023) with a different choice of hyperparameters. Standard empirical risk minimization (ERM) for MLPs achieves nearly optimal accuracy and multicalibration error.
@@ 1.1 Our Contributions
We find that in practice, this comparison is surprisingly subtle: multicalibration algorithms do not always improve multicalibration (relative to the ERM baseline), for example. From the results of our extensive experiments on tabular, vision, and language tasks (involving running multicalibration algorithms more than 45K times), we extract a number of observations clarifying the utility of multicalibration. Most significantly, we find:
@@
    Multicalibration algorithms are very sensitive to hyperparameter choices, and can require large parameter sweeps to avoid overfitting. These algorithms tend to be most effective in regimes with large amounts of available data, such as image and language datasets.
@@
    Traditional calibration methods such as Platt scaling or isotonic regression can sometimes give nearly the same performance as multicalibration algorithms, and are hyperparameter-free. Furthermore, compared to multicalibration post-processing, they are extremely computationally efficient and do not take very long to evaluate.
We also present numerous practical takeaways for users of multicalibration algorithms, which are not apparent from the existing theoretical literature, but are crucial considerations in practice. We believe that our investigations will not only broaden the practical applicability of multicalibration as a concept and algorithm, but also provide valuable information to the theoretical community as to what barriers multicalibration faces in practice. To both of these ends, all code used in our experiments is publicly accessible [^1] https://github.com/dutchhansen/multicalibration.
Organization. In Section 1.2, we begin with a brief review of related theoretical and experimental work in the multicalibration literature. We then detail our key experimental design choices in Section 2, before discussing our results on tabular data in Section 3. We extend our results to more complex image and language datasets in Section 4. Finally, we conclude with limitations of our experiments as well as practical takeaways for practitioners of fair machine learning in Section 5.
@@ 1.2 Related Works: Theory and Practice
There is reason to believe that empirical risk minimization (ERM) on neural networks and other machine learning models may result in multicalibrated predictors. In recent works, Błasiok et al., (2024, 2023) prove that loss minimization with neural networks may yield multicalibrated predictors. Their proofs, however, may not be directly applicable to practice as they rely on an idealized optimization procedure. Nonetheless, both works echo a relationship between ERM and multicalibration also articulated in Liu et al., (2019), who show that group-wise calibration may be an inevitable consequence of well-performing models.
@@
In recent work, Detommaso et al., (2024) utilize multicalibration as a tool to improve the overall uncertainty and confidence calibration of language model but, to our knowledge, do not focus on or report fairness towards protected subgroups. Like us, they point out various issues with the standard multicalibration algorithm, which they address with early stopping and adaptive binning. We instead perform a large hyperparameter sweep which effectively implements an early stopping mechanism. We discuss this further in Section A.1. Nonetheless, our results for large models are complementary to those of Detommaso et al., (2024): both works demonstrate that (1) standard multicalibration can at times be difficult to get working in practice; and (2) ideas from the theoretical multicalibration literature can have impact at the scale of large models. We provide additional discussion of related works in model calibration and subgroup robustness in Appendix A.
@@ 2 Preliminaries
We work in the binary classification setting with a domain $\mathcal{X}$ and binary label set $\mathcal{Y}=\{0,1\}$, and assume data is drawn from a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$. We consider arbitrary risk predictors $f:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$, which return probability distributions over the binary label space. We will measure the calibration of $f$ on a dataset $(X,Y)\in(\mathcal{X}\times\mathcal{Y})^{n}$ with the binned variant of the well-known and standard Expected Calibration Error, which we refer to as ECE (Guo et al., , 2017). Throughout, we measure ECE with 10 bins of equal width $0.1$.
Recent work has questioned ECE as a calibration measure, due to consistency and continuity issues that come with relying on a fixed bin width (Błasiok et al., , 2023). To address these, we also report calibration as measured by smoothed ECE (smECE) (Błasiok and Nakkiran, , 2023), which (1) can be roughly thought of as the ECE after applying a suitable kernel-smoothing to the predictions, and; (2) satisfies desirable continuity and consistency guarantees (Błasiok et al., , 2023; Błasiok and Nakkiran, , 2023). Importantly, unlike binned ECE, there are no hyperparameters associated with measuring the smoothed calibration error. A full description of smECE is beyond the scope of our work—we refer the interested reader to Błasiok and Nakkiran, (2023).
Multicalibration requires that a predictor have not only small calibration error overall, but also when restricted to marginal subgroup distributions of the data. In particular, we assume that there is a (finite) collection of groups $G=\{g_{1},g_{2},\dots\}$, where $g_{i}\subseteq\mathcal{X}$. We operationalize measuring multicalibration by reporting the maximum calibration error over a given collection of subgroups $G$. Taking the max avoids fairness concerns associated with the (weighted) mean of groups of varying size and/or degree of overlap. Note that the subgroup collection $G$ is context and dataset dependent, and that the groups within $G$ may be overlapping, capturing desirable intersectionality notions (Ovalle et al., , 2023).
@@ 2.1 Multicalibration Post-Processing Algorithms and Hyperparameter Selection
The recent work of Haghtalab et al., (2023) also provide a family of alternative multicalibration algorithms with better theoretical sample complexity guarantees. This is motivated by the fact that HKRR is known to be theoretically sample inefficient (Gopalan et al., 2022b, ), and easily overfits (Detommaso et al., , 2024). At a high level, each algorithm of Haghtalab et al., (2023) corresponds to a certain two-player game. Different algorithms in the family are a consequence of each player playing a different online learning algorithm. We detail the hyperparameters we over which we search in Section D.2, but note here that we use the same code and predominantly the same parameters reported by the authors. We refer to any (post-processing) algorithm in this family as HJZ.
@@
In addition to these algorithms, we test the usefulness of three standard calibration techniques in reducing multicalibration error: Platt scaling (Platt et al., , 1999), isotonic regression (Zadrozny and Elkan, , 2002), and temperature scaling Guo et al., (2017). The first two techniques are hyperparameter-free, and we use implementations given by Scikit-learn. We also use a parameter-free version of temperature scaling which we detail in Section D.3.
2.2 Experimental Methodology
@@
Our investigation is primarily motivated by fairness desiderata. Therefore, we examine datasets where features or metadata allow us to naturally define meaningful collections of protected groups $G$. This includes five tabular datasets (ACS Income, UCI Bank Marketing, UCI Credit Default, HMDA, MEPS), two language datasets (Civil Comments, Amazon Polarity), and two image datasets (CelebA, Camelyon17). We detail and provide citations for each of our datasets and subgroups in Appendix C. For each dataset, we define between 10 and 20 overlapping subgroups depending on available features or metadata. For the tabular datasets, these subgroups are defined by features and conjunctions of features. For the image datasets, we use relevant metadata not available to the model in order to define subgroups. For the language datasets, we use the presence of meaningful words.
@@
For consistency, we partition all datasets into three subsets: training, validation, and test. Test sets remain fixed across all experiments. We report accuracy and multicalibration metrics on the test set averaged over five splits of train and validation sets for tabular data, and three splits for more complex data. In each experiment where a multicalibration post-processing algorithm is used, we draw a holdout of variable size from our training set, which we term the calibration set. We term the fraction of our training set used in post-processing the calibration fraction. The calibration fractions over which we search appear in Section E.1 for tabular datasets and Section E.2 for image and language datasets.
The calibration set is used solely in multicalibration post-processing, and is not used in training a model prior to the post-process. This procedure is motivated by a need to measure the importance of fresh samples in multicalibration post-processing. If a model is already multicalibrated on its entire training set $S$, we cannot re-use $S$ in HKRR or HJZ to improve the model, since the algorithms cannot improve on a predictor which is already perfectly multicalibrated on a particular dataset. This applies to models such as, for example, neural networks, which usually fit their training set to very low calibration error and high accuracy (Carrell et al., , 2022). For these models, we also anticipate that the multicalibration error on the training set will be low, and hence, the data from $S$ unusable for post-processing. Therefore, the calibration fraction itself is an important hyperparameter we consider. Ideally, in order to maximize the resulting accuracy of the final model, we would utilize as much data as possible for model training, and minimize the amount of data required for multicalibration post-processing. However, due to the sample complexity of multicalibration algorithms, we will see that finding this specific point can be difficult (see Figure 3).

### v2

@@ Abstract
Calibration is a well-studied property of predictors which guarantees meaningful uncertainty estimates. Multicalibration is a related notion — originating in algorithmic fairness — which requires predictors to be simultaneously calibrated over a potentially complex and overlapping collection of protected subpopulations (such as groups defined by ethnicity, race, or income). We conduct the first comprehensive study evaluating the usefulness of multicalibration post-processing across a broad set of tabular, image, and language datasets for models spanning from simple decision trees to 90 million parameter fine-tuned LLMs. Our findings can be summarized as follows: (1) models which are calibrated out of the box tend to be relatively multicalibrated without any additional post-processing; (2) multicalibration post-processing can help inherently uncalibrated models and large vision and language models; and (3) traditional calibration measures may sometimes provide multicalibration implicitly. More generally, we distill many independent observations which may be useful for practical and effective applications of multicalibration post-processing in real-world contexts. We also release a python package implementing multicalibration algorithms, available via ‘pip install multicalibration’.
@@ 1 Introduction
Although there is a host of theoretical results surrounding multicalibration and related notions, there is little systematic empirical study of the latent multicalibration error of popular machine learning models, the effectiveness of multicalibration post-processing algorithms, or even best practices for practitioners who wish to apply ideas and algorithms from the multicalibration literature. In particular, theoretical results are often concerned with multicalibration towards subgroups defined by potentially infinite hypothesis classes (Hébert-Johnson et al., , 2018; Haghtalab et al., , 2023). In contrast, fairness practitioners may prioritize the equitable performance of a model over a finite number of protected subgroups of interest. These groups are typically defined by attributes and meta-data such as race, sex, ethnicity, etc. (Chen et al., , 2023) which are normatively deemed as important. Furthermore, most existing works applying multicalibration in practical settings only focus on one-off datasets or examples, and do not validate the algorithm(s) across a variety of datasets and models or with realistic finite sample restrictions (Liu et al., , 2019; Barda et al., , 2020; La Cava et al., , 2023).
To address these, we consider a “realistic” setup where a practitioner only has a finite amount of data, and must choose how to partition this data between learning and post-processing in order to achieve a suitable accuracy and multicalibration error rate over a finite set of subgroups. This allows us to investigate many important questions pertaining to the practical usage of multicalibration concepts and algorithms, which, to the best of our knowledge, have not been systematically considered by the theoretical or practical communities. For example, we use this setup to investigate the effectiveness of multicalibration post-processing algorithms and hyperparameter choices, as well as the latent multicalibration properties of popular machine learning models at a large scale.
More broadly, we initiate a systematic empirical study of multicalibration with the goal of answering two salient questions:
@@
The conventional wisdom is that multicalibration is something that is not naturally achieved by ML algorithms—this is precisely why many in the community have focused on creating post-processing algorithms which do achieve it (see e.g. Hébert-Johnson et al., (2018); Gopalan et al., 2022b , and Section 1.2). However, recent theoretical results suggest that multicalibration may in fact be an inevitable consequence of certain empirical risk minimization (ERM) methods with proper losses (Błasiok et al., , 2023; Liu et al., , 2019). This apparent conflict between conventional wisdom and recent results has not been tested in practice. We propose studying Question 1 since we believe that the current state of multicalibration in ML models should be systematically studied to better understand the implications for modern learning setups involving large models and fine-tuning. Question 2 is complementary and focused on investigating the effectiveness of current multicalibration algorithms on real datasets and illuminating challenges which can guide the development of future algorithms. A partial answer to one or both of these questions could help practitioners concerned about fairness understand when they should or should not expect multicalibration algorithms to help.
Refer to captionRefer to captionRefer to caption
Figure 1: Test accuracy vs. maximum group-wise calibration error (smECE) averaged over five train/validation splits for simple neural networks (MLPs) trained on Credit Default, MEPS, and ACS Income. Each point corresponds to the performance of the multicalibration post-processing algorithm HKRR (Hébert-Johnson et al., , 2018) or HJZ (Haghtalab et al., , 2023) with a different choice of hyperparameters. Standard empirical risk minimization (ERM) for MLPs achieves nearly optimal accuracy and multicalibration error. Full hyperparameter plots for each base dataset are in Appendix F.
@@ 1.1 Our Contributions
We find that in practice, this comparison is surprisingly subtle: multicalibration algorithms do not always improve worst group calibration error (relative to the ERM baseline), for example. From the results of our extensive experiments on tabular, vision, and language tasks (involving running multicalibration algorithms more than 45K times), we extract a number of observations clarifying the utility of multicalibration algorithms. Most significantly, we find:
@@
    Multicalibration algorithms are very sensitive to hyperparameter choices, and can require large parameter sweeps to avoid overfitting. Furthermore, these algorithms tend to be most effective in regimes with large amounts of available data, such as image and language datasets.
@@
    Traditional calibration methods such as Platt scaling or isotonic regression can sometimes give nearly the same performance as multicalibration algorithms, and are hyperparameter-free. Furthermore, compared to multicalibration post-processing, they are extremely computationally efficient.
We also present numerous practical takeaways for users of multicalibration algorithms, which are not apparent from the existing theoretical literature, but are crucial considerations in practice. We believe that our investigations will not only broaden the practical applicability of multicalibration as a concept and algorithm, but also provide valuable information to the theoretical community as to what barriers multicalibration faces in practice. To both of these ends, all code used in our experiments is publicly accessible, and we also release a python package implementing two multicalibration algorithms which we make available via ‘pip install multicalibration’.  [^1] Code for our experiments is available at https://github.com/dutchhansen/empirical-multicalibration, while code for the python package is available at https://github.com/sid-devic/multicalibration.
Organization. In Section 1.2, we begin with a brief review of related theoretical and experimental work in the multicalibration literature. We then detail our key experimental design choices in Section 2, before discussing our results on tabular data in Section 3. We extend our results to more complex image and language datasets in Section 4. In Section 5, we investigate to what extent multicalibration post-processing can re-use data from model training. Finally, we conclude with limitations of our experiments as well as practical takeaways for practitioners of fair machine learning in Section 6.
@@ 1.2 Related Works: Theory and Practice
There is reason to believe that empirical risk minimization (ERM) on neural networks and other machine learning models may result in multicalibrated predictors. In recent works, Błasiok et al., (2024, 2023) prove that loss minimization with neural networks may yield multicalibrated predictors. Their proofs, however, may not be directly applicable to practice as they rely on an idealized optimization procedure (we provide further discussion of the relation between our works in Section A.1). Nonetheless, both works echo a relationship between ERM and multicalibration also articulated in Liu et al., (2019), who show that group-wise calibration may be an inevitable consequence of well-performing models.
@@
In recent work, Detommaso et al., (2024) utilize multicalibration as a tool to improve the overall uncertainty and confidence calibration of language models but, to our knowledge, do not focus on or report fairness towards protected subgroups. Like us, they point out various issues with the standard multicalibration algorithm, which they address with early stopping and adaptive binning. We instead perform a large hyperparameter sweep which effectively implements an early stopping mechanism. We discuss this further in Section A.2. Nonetheless, our results for large models are complementary to those of Detommaso et al., (2024): both works demonstrate that (1) standard multicalibration can at times be difficult to get working in practice; and (2) ideas from the theoretical multicalibration literature can have impact at the scale of large models. We provide additional discussion of related works in model calibration and subgroup robustness in Appendix A.
@@ 2 Preliminaries
We work in the binary classification setting with a domain $\mathcal{X}$ and binary label set $\mathcal{Y}=\{0,1\}$, and assume data is drawn from a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$. We consider arbitrary risk predictors $f:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$, which return probability distributions over the binary label space. We will measure the calibration of $f$ on a dataset $S\in(\mathcal{X}\times\mathcal{Y})^{n}$ with the binned variant of the well-known and standard Expected Calibration Error, which we refer to as ECE (Guo et al., , 2017). Throughout, we measure ECE with 10 bins of equal width $0.1$.
Recent work has questioned ECE as a calibration measure, due to consistency and continuity issues that come with relying on a fixed bin width (Błasiok et al., , 2023). To address these, we also report calibration as measured by smoothed ECE (smECE) (Błasiok and Nakkiran, , 2023), which (1) can be roughly thought of as the ECE after applying a suitable kernel-smoothing to the predictions, and (2) satisfies desirable continuity and consistency guarantees (Błasiok et al., , 2023; Błasiok and Nakkiran, , 2023). Importantly, unlike binned ECE, there are no hyperparameters associated with measuring the smoothed calibration error. A full description of smECE is beyond the scope of our work—we refer the interested reader to Błasiok and Nakkiran, (2023).
Multicalibration requires that a predictor have not only small calibration error overall, but also when restricted to marginal subgroup distributions of the data. In particular, we assume that there is a (finite) collection of groups $G=\{g_{1},g_{2},\dots\}$, where $g_{i}\subseteq\mathcal{X}$. We operationalize measuring multicalibration by reporting the maximum calibration error over a given collection of subgroups $G$ [^2] Multicalibration was introduced before smECE, and was designed to reduce a bucketed group-wise calibration error (similar to ECE). Therefore, our investigations concerning smECE are of a purely empirical character. Taking the max avoids fairness concerns associated with the (weighted) mean of groups of varying size and/or degree of overlap. Note that the subgroup collection $G$ is context and dataset dependent, and that the groups within $G$ may be overlapping, capturing desirable intersectionality notions (Ovalle et al., , 2023).
@@ 2.1 Multicalibration Post-Processing Algorithms and Hyperparameter Selection
The recent work of Haghtalab et al., (2023) also provide a family of alternative multicalibration algorithms with better theoretical sample complexity guarantees. This is motivated by the fact that HKRR is known to be theoretically sample inefficient (Gopalan et al., 2022b, ), and easily overfits (Detommaso et al., , 2024). For example, the number of samples required for generalization guarantees of HKRR is typically $O(\tfrac{1}{\alpha^{4}\lambda^{1.5}})$, where $\alpha$ determines the allowed multicalibration violation and $\lambda$ represents a suitable discretization width. For reasonably small values of $\alpha$ and $\lambda$, this can balloon the required number of samples to an unreasonable number. At a high level, each algorithm of Haghtalab et al., (2023) corresponds to a certain two-player game. Different algorithms in the family are a consequence of each player playing a different online learning algorithm. We detail the hyperparameters over which we search in Section D.2, but note here that we use the same code and predominantly the same parameters reported by the authors. We refer to any (post-processing) algorithm in this family as HJZ.
@@
In addition to the multicalibration algorithms HJZ and HKRR, we test the usefulness of three standard calibration techniques in reducing multicalibration error: Platt scaling (Platt et al., , 1999), isotonic regression (Zadrozny and Elkan, , 2002), and temperature scaling Guo et al., (2017). The first two techniques are hyperparameter-free, and we use implementations given by Scikit-learn. We also use a parameter-free version of temperature scaling which we detail in Section D.3.
2.2 Subgroup Selection and Experimental Methodology
Multicalibration post-processing requires the selection of “groups” or subsets of the population of interest. As our investigation is primarily motivated by fairness desiderata, these subgroups determine what segments of the population the practitioner would like to “protect” or guarantee performance over. In most practical applications, these subgroups are constructed via features or conjunctions of features given as input for each data point. This way of constructing groups is standard: it is used by large production systems such as LinkedIn (Quiñonero Candela et al., , 2023), in the measurement of bias in ML (Tifrea et al., , 2024; Atwood et al., , 2024) and NLP systems (Baldini et al., , 2022; Li et al., , 2023), and in the auditing of large, deployed ML systems like Facebook (Ali et al., , 2019; Imana et al., , 2024). Furthermore, this approach still takes advantage of group intersectionality notions (Ovalle et al., , 2023), a cornerstone of multigroup fairness.
Remark 1
Using conjunctions of features or additional meta-data is not the only way to construct subgroups of the data. For any imperfect predictor $\hat{p}$, we can (nearly) always construct groups against which the predictor is not multicalibrated. For example, simply take the set of data points for which $\hat{p}$ predicts the incorrect label. Groups defined in this way may potentially be “as complex” as the underlying predictor $\hat{p}$. Nonetheless, it is not clear whether this is a meaningful set of groups to ask for multicalibration against (as it may require post-processing $\hat{p}$ to be a perfect predictor). To avoid such discussion, we intentionally determine groups by available features which we hypothesize practitioners may deem important or “sensitive” to the underlying prediction task.
@@ 2.2 Subgroup Selection and Experimental Methodology
We experiment across a variety of classification tasks: five tabular datasets (ACS Income, UCI Bank Marketing, UCI Credit Default, HMDA, MEPS), two language datasets (Civil Comments, Amazon Polarity), and two image datasets (CelebA, Camelyon17). For each dataset, we also define between 10 and 20 overlapping subgroups depending on available features or metadata. We detail and provide citations for each of our datasets and exact subgroup descriptions in Appendix C. In what follows, we give a high level overview of how we determined subgroups in our experiments.
For our tested tabular datasets, we determined groups by “sensitive” attributes—individual characteristics against which practitioners would not want to discriminate. In many cases, such attributes naturally include race, gender, and age, and vary with available information. For example, on ACS Income, we include groups such as “Multiracial,” “American Indian,” and “Seniors.” We also include some groups which are conjunctions of two attributes, for example “Black Women” or “White Women.” On Credit Default, for example, we include groups such as “Education = High School, Age > 40,” or “Married, Age > 60.” Similar groups are used throughout all tabular datasets, as well as CelebA.
On datasets where samples are not in correspondence with individuals—Camelyon17, Amazon Polarity, and Civil Comments—we define groups based on available information that can be viewed as “sensitive” with respect to the underlying task. In other words, we define groups such that an individual or institution using a predictor which is miscalibrated on this group may be seen as discriminating against the group. For example, a social media service should ideally not be underconfident when predicting the toxicity of posts mentioning a minority identity group; such predictions may allow hate speech to remain on the platform, or may provide differential engagement boosting based on the presence of racial identifiers in posts. Therefore, we include “Muslim,” “Black,” and various other phrases defining protected groups in the Civil Comments dataset. In the same vein, one would want a shopping platform to promote product listings fairly: a predictor which outputs mis-multicalibrated product ratings may boost listings for technology products proportionally to their true ratings but unfairly ignore the positive reviews of book listings. Thus, some groups from the Amazon Polarity dataset are defined as reviews containing the phrases “book,” “food,” “exercise,” or “tech.”
In Section 6.1, we further discuss group selection methodology and speculate about other ways of achieving multicalibration via group design.
@@
For consistency, we partition all datasets into three subsets: training, validation, and test. Test sets remain fixed across all experiments. We report accuracy and multicalibration metrics on the test set averaged over five random splits of train and validation sets for tabular data, and three splits for more complex data. Whenever a (multi)calibration post-processing algorithm is used, we run it using a holdout set of variable size from our training set, which we term the calibration set. We define the fraction of the training set used in (multi)calibration post-processing to be the calibration fraction. The exact calibration fractions over which we search appear in Section E.1 for tabular datasets and Section E.2 for image and language datasets. Note that multicalibration post-processing methods are far less sample efficient than standard post-hoc calibration methods. Therefore, the calibration fractions we test are broadly distributed between 5% and 100% of the training data (rather than using, say, a standard 10% of data for post-hoc calibration).
The calibration set is used solely in multicalibration post-processing, and is not used in training a model prior to the post-process. This procedure is motivated by a need to measure the importance of fresh samples in multicalibration post-processing. If a model is already multicalibrated on its entire training set $S$, we cannot re-use $S$ in HKRR or HJZ to improve the model, since the algorithms cannot improve on a predictor which is already perfectly multicalibrated on a particular dataset. This applies to models such as neural networks, which usually fit their training set to very low calibration error and high accuracy (Carrell et al., , 2022). For these models, we also anticipate that the multicalibration error on the training set will be low, and hence, the data from $S$ unusable for post-processing. [^3] Indeed, we test this more rigorously in Section 5, where we experiment with data reuse between model training and multicalibration post-processing. Therefore, the calibration fraction itself is an important hyperparameter we consider. Ideally, in order to maximize the resulting accuracy of the final model, we would utilize as much data as possible for model training, and minimize the amount of data required for multicalibration post-processing. However, due to the sample complexity of multicalibration algorithms, we will see that finding this specific point can be difficult (see Figure 3).